# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v2

@@ 3.2 Bandit Algorithms for Adaptive RM Selection
We employ LinUCB (Li et al., 2010), a contextual bandit algorithm for the arm or RM selection. We choose LinUCB because it is a contextual bandit algorithm that effectively leverages context information, making it well-suited for integration into LLM frameworks due to its ability to dynamically adapt decisions based on contextual embeddings. Additionally, LinUCB has been shown to provide a good trade-off between computational efficiency and performance (Zhou, 2015). LinUCB assumes that the MAB reward can be modeled linearly as a function of context features and computes the expected MAB reward of each arm with an upper confidence bound to ensure exploration (Garivier and Moulines, 2008, 2011). In each step $t$, we have a batch of input prompts $\mathbf{x}_{m,t}$ for which we compute sentence embeddings, using the policy model $\pi_{m}$, and use the mean sentence embedding as the context $c(t)$ to the MAB, i.e., $c(t)=\sum_{x\in\mathbf{x}_{m,t}}e_{m}(x)/|\mathbf{x}_{m,t}|$, where $e_{m}(.)\in\mathbb{R}^{d}$ yields the sentence embedding from the model $\pi_{m}$. We calculate the embedding for a prompt as the last token embedding from model $\pi_{m}$ (details in Sec. A.1). The learned parameters of the LinUCB bandit include $\hat{\theta}_{k}\in\mathbb{R}^{d}$ which represents the learned weights for the features of each reward model and $A_{k}\in\mathbb{R}^{d\times d}$ (a covariance matrix) and a bias vector $b_{k}\in\mathbb{R}^{d}$ corresponding to each arm or RM $R_{k}$. We initialize the parameters for LinUCB by randomly initializing $b_{k}$ and setting parameter $A_{k}$ to the identity matrix. Based on the LinUCB algorithm, for each batch, the selected RM $R^{\star}_{t}$ is determined by $R^{\star}_{t}=R_{j}$ such that:
@@
where $\hat{\theta}_{k}=A_{k}^{-1}b_{k}$. $A_{k}$ and $b_{k}$ are updated based on the MAB reward for each RM, which corresponds to the normalized negative train loss ^$-\hat{\mathcal{L}}^{m}$ (described in detail in Sec. A.2):
@@
We emphasize that the effectiveness of our approach is not tied to particular bandit algorithm but rather from the exploration-exploitation trade-off of MAB framework. LASeR is agnostic to the bandit algorithms and can incorporate alternatives to LinUCB. Indeed, we provide a comparison of LinUCB and other bandit algorithms in Appendix C.
@@ 3.3 LLM and Bandit Training in LASeR
A key aspect of our approach is the generation of new preference training data in each iteration using the generations of the LLM itself and the RM selected by the MAB. Fig. 1 presents our training procedure, broken down into three stages: (i) the MAB selects an RM $R^{\star}_{t}$ (see Sec. 3.2; Fig. 1 left), generating preference pairs by scoring the LLM’s outputs using the RM (Fig. 1 (top-right)), and parameter updates to the LLM and MAB. In this way, the model continuously learns from its own outputs, guided by the selected reward model. After each LLM train step (i.e., one mini-batch), the MAB’s parameters are updated based on the observed MAB reward, i.e., how much the LLM’s loss decreased from using the selected RM. In the case of LinUCB, this involves updating the parameter estimates $b_{k},A_{k}$ (see Fig. 1; bottom in green). This entire process – selection of reward models, generation of new supervision data, fine-tuning, and bandit updates – repeats for a total of $M$ iterations (summarized in Algorithm 1).
LASeR with Best-of-$n$ Sampling.   For settings where finetuning the LLM is not desirable or feasible, LASeR can also be applied to learn the MAB parameters without training the LLM. Rather than fine-tuning the model with preference data, we employ best-of-$n$ sampling (Lightman et al., 2023; Sun et al., 2024), where multiple responses are generated, and the best one is selected based on the RM. The bandit parameters are then updated using equation (2), with the MAB reward calculated as the negative normalized NLL loss on the train data. This updated bandit can subsequently be used for inference on the test set. This approach is particularly useful for long-context understanding tasks, where training would be too computationally intensive (see setting in Sec. 4.2).
@@ 4.1 Experimental Setup
Models.   We conduct our experiments on the Llama-3-8B base (AI@Meta, 2024) model. We present additional results with Mistral-7b-v3-Instruct (Jiang et al., 2023) and Qwen2.5-32B (Qwen Team, 2024a) in Appendix B. For training, all models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficiency. For both training and inference, we do 0-shot prompting and sample $n\!=\!30$ responses per prompt with temperature $0.8$ (see Sec. A.1 for more details).
Reward Models.   We select $K=4$ strong 7B RMs from RewardBench (Lambert et al., 2024), which include Zephyr-7B-Alpha, Qwen1.5-7B-Chat, Eurus-7B-KTO, and OLMo-7B-Instruct. Following the pipeline outlined in Lambert et al. (2024), for these models, we compute the reward for each response as the log-likelihood of the RM for that response (see Secs. A.1 and D for more details and discussion about the RM choices).
Datasets and Metrics.   Our experiments cover a range of tasks and datasets (see Sec. A.1):
@@
  • Long-Context Understanding: As finetuning LLMs on long-context inputs is computationally intensive, we demonstrate the effectiveness of LASeR using Best-of-$n$ sampling on LongBench (Bai et al., 2023), which consists of multiple tasks, such as single-document QA, multi-document QA, summarization, and few-shot learning. For summarization, we report Rouge-L (Lin, 2004), whereas for the rest, we report the F1 score.
@@
  • Best RM: From our collection of RMs, we pick the RM that corresponds to the best overall score on RewardBench (Lambert et al., 2024): Zephyr-7B-Alpha. We use this single RM during training (c.f. Sec. 3.1). This baseline reflects the performance gain a user could expect when selecting the best RM from a leaderboard without knowing a priori how it generalizes to a particular task.
@@
  • Classifier Selection: To compare against a context-sensitive baseline that does not use an MAB, we train a $K$-way classifier to perform RM selection using data from RewardBench (see Sec. A.1). Specifically, for each query and RM, we compute the RM’s score of the annotated preferred and dispreferred response. The RM that assigns the correct preference ordering with the highest difference between the scores of the preferred and dispreferred responses is chosen as the RM label and used to train the classifier. While training the LLM, for each input $x_{i}\in\mathcal{D}$, we select the RM for building preference pairs based on this trained classifier.
@@
  • RM Agreement Ensemble: Because ensembling scores through averaging is sensitive to the absolute scores produced (which may differ between RMs), we follow Wang et al. (2024d) in ensembling through ranking and agreement. Specifically, we generate 32 responses for each query and sample 100 pairs for each. We score each pair with RMs, constructing a preference dataset by choosing the 10 pairs with the highest agreement of preference rankings across RMs.
@@
Conceptually, among the single RM selection methods, the best RM baseline serves as an “exploit-only” setting that exploits the best available RM based on aggregate RewardBench scores. On the other hand, the random and sequential selection baselines are “explore-only” in that they pick a new RM either randomly or via a predefined sequence, irrespective of the performance of each arm (RM). LASeR and the classifier approach represent two methods for selecting a single optimal RM for each instance; however, we note that the classifier approach depends on a fine-grained, in-distribution dataset of queries paired with their corresponding suitable RMs for annotating preferences, which is often absent in practice. In our experiments, we demonstrate the benefits of balancing the exploration-exploitation trade-off compared to “exploit-only” and “explore-only” approaches. Additionally, different RM ensemble baselines employ various methods for combining RMs while utilizing them simultaneously. Furthermore, we analyze potential conflicting signals among RMs to highlight the advantages of our framework over RM ensemble baselines (Sec. 5). The settings for training LASeR and baselines are provided in Sec. A.1.
@@ 4.2 Main Results
LASeR achieves the best average accuracy on reasoning tasks.   Table 1 shows our method consistently outperforms the baselines across multiple benchmarks, particularly in the StrategyQA and GSM8K tasks. For example, LASeR improves approximately $2\%$ absolute accuracy over the sequential baseline on both GSM8K and StrategyQA. In cases where the best RM is not known beforehand, LASeR surpasses the performance of the average RM by $2.7\%$ and the RM Score Ensemble for each instance by $2.67\%$ (in accuracy averaged over the three tasks). Moreover, this lower performance by the RM Score baseline is not purely due to variance in the scores: LASeR also surpasses the RM Agreement Ensemble and RM Online Ensemble by $0.91\%$ and $2.27\%$, respectively. Overall, LASeR provides consistent results while the underlined second-place models show inconsistent performance across datasets. These results also emphasize the benefit of LASeR, as it eliminates the need to choose a different RM in advance or ensemble multiple RMs.
LASeR beats baselines at instruction-following.   Often, LLMs are used by large numbers of people with a diverse set of queries, goals, and intentions, and their preferences vary based on the underlying query. To demonstrate the effectiveness of LASeR in such settings,
@@
we compare the instruction-following performance in Fig. 2, i.e., AlpacaEval win rates, of LLMs trained using LASeR with the baselines using WildChat. Specifically, LASeR achieves substantial win rates compared to single RM selection baselines such as random and sequential selection, with $78.33\%$, and $71.45\%$, respectively. We also outperform training with the single best RM (per RewardBench) by a $56.34\%$ win rate. Compared to RM ensemble baselines, LASeR achieves $72.69\%$ and $67.88\%$ win rates over RM Score Ensemble and RM Online Ensemble, respectively. We hypothesize the lower win rate of the baselines stems from the inability of these baselines to deal with conflicting signals from multiple RMs (see Fig. 3 for further analysis). Overall, these results highlight that LASeR excels without gold labels and performs consistently well at following instructions across various user queries, showcasing its adaptability to diverse tasks.
LASeR’s adaptive RM selection helps long-context understanding.   Given the cost of training long-context systems, for LongBench (Bai et al., 2023), rather than finetuning a model using RMs, we employ the selected RM to rerank generation in Best-of-$n$ sampling (see Sec. 3.3).
@@
In Table 2, we observe that LASeR consistently outperforms the baselines across tasks. LASeR improves single-doc QA (SiQA) by $3.58$ F1 points over the base Llama-3-8B model and $2.64$ F1 points over random RM selection. On multi-doc QA (MuQA), our approach improves performance over the Llama-3-8B base model by $\approx\!4$ F1 points. Furthermore, on few-shot (FS) learning tasks, LASeR provides over $3$ points gain in F1 compared to the base model, surpassing the average RM performance by up to $2.4$ F1 points and demonstrating its effectiveness across tasks. Lastly, Table 2 demonstrates that LASeR consistently outperforms the RM Score Ensemble baseline across different long-context tasks (except for the Summarization task where LASeR is comparable to Best RM), e.g., a $\approx\!3$ F1 point boost on single-doc QA and few-shot learning tasks.
@@ 5 Additional Analysis of LASeR
Presence of Conflicting Signals among RMs.   In Sec. 4.2, we find that LASeR consistently outperforms other RM ensemble baselines across a wide variety of tasks. We attribute some of these performance gains to the inability of the multi-RM baseline to handle
@@
conflicting signals, resulting in subpar training data from multiple RMs. To study this, we sample pairs of outputs generated by Llama-3-8B on MMLU as well as WildChat and evaluate the consistency of response preferences measured by multiple RMs. Since pair-wise preferences are binary, we compute F1 to measure consistency with one RM’s preferences serving as the reference. Fig. 3 (left on MMLU) reveals that Qwen and Zephyr have the highest agreement rate at $0.77$, while Qwen’s agreement with Eurus and Olmo is lower at $0.58$ and $0.43$, respectively. This is expected as Qwen and Zephyr are the top-performing models in reasoning according to RewardBench, while Olmo ranks the lowest in reasoning ability among the four models. We observe similar trends in agreement across RMs on WildChat (albeit with different agreement scores), which contains user queries asked LLMs in the wild; see Fig. 3 (right). It appears that for more heterogeneous datasets with more categories, the level of disagreement among RMs (or conflict) increases. This also highlights LASeR’s advantages over multi-RM baselines that do not address conflicts in RMs and may explain why choosing one RM in LASeR and the best RM baseline outperforms multi-RM ensembles.
@@
Training Efficiency of LASeR.
We show the accuracy-training time tradeoff in Fig. 4 by comparing the GSM8K performance of training with LASeR and different baselines, along with the corresponding wall clock training time. Wall clock time is measured as the training time of a model (hours), keeping compute resource consistent. We find that sequentially optimizing over each RM performs the worst in terms of training time (3×3$\times$3 × of LASeR) while RM score ensemble has the worst accuracy (and takes 2×2$\times$2 × the training time of LASeR). Moreover, LASeR outperforms all other baselines in terms of accuracy while maintaining the lowest training time, being more than twice as fast as the second-best baseline. LASeR’s efficiency comes from fast convergence compared to the sequential selection baseline and avoiding the overhead of loading and evaluating multiple RMs per step, as required in RM ensemble methods (Online Ensemble, Score Ensemble, and RM Agreement Ensemble). Instead, it selects and uses only one RM per mini-batch, saving GPU memory and compute. Ensemble baselines require scoring all candidate responses with all RMs, whereas LASeR scores responses with only the selected RM, reducing overhead. Lastly, we note that LASeR can be run in an offline data creation mode, where RM selection and scoring are done once to generate preference data, unlike online ensemble methods that require multiple RM passes per step.

### v3

@@ 3.2 Bandit Algorithms for Adaptive RM Selection
We employ LinUCB (Li et al., 2010), a contextual bandit algorithm for the arm or RM selection. We choose LinUCB because it is a contextual bandit algorithm that effectively leverages context information, making it well-suited for integration into LLM frameworks due to its ability to dynamically adapt decisions based on contextual embeddings. Additionally, LinUCB has been shown to provide a good trade-off between computational efficiency and performance (Zhou, 2015). Therefore, it can be incorporated into any iterative LLM training framework with minimal overhead (described in detail in Sec.˜A.1). LinUCB assumes that the MAB reward can be modeled linearly as a function of context features and computes the expected MAB reward of each arm with an upper confidence bound to ensure exploration (Garivier and Moulines, 2008, 2011). In each step $t$, we have a batch of input prompts $\mathbf{x}_{m,t}$ for which we compute sentence embeddings, using the policy model $\pi_{m}$, and use the mean sentence embedding as the context $c(t)$ to the MAB, i.e., $c(t)=\sum_{x\in\mathbf{x}_{m,t}}e_{m}(x)/|\mathbf{x}_{m,t}|$, where $e_{m}(.)\in\mathbb{R}^{d}$ yields the sentence embedding from the model $\pi_{m}$. We calculate the embedding for a prompt as the last token embedding from model $\pi_{m}$. We use the last-token embedding as the context representation because in Transformer-based LLMs, this position typically aggregates information from the entire sequence. Moreover, this method is a standard and commonly used approach for extracting LLM embeddings (Wang et al., 2024d; BehnamGhader et al., 2024). We provide more details regarding token embedding extraction in Sec.˜A.1 and an ablation study comparing context embedding methods in Appendix˜D. Even if individual batches vary, the repeated exposure to diverse inputs allows the bandit to learn which RMs are more helpful overall. Over training iterations, this appears to provide a robust signal for the bandit to learn effective RM selection, as demonstrated by LASeR’s consistent gains across datasets and the convergence behavior of MAB rewards shown in Appendix˜D.
The learned parameters of the LinUCB bandit include $\hat{\theta}_{k}\in\mathbb{R}^{d}$ which represents the learned weights for the features of each reward model and $A_{k}\in\mathbb{R}^{d\times d}$ (a covariance matrix) and a bias vector $b_{k}\in\mathbb{R}^{d}$ corresponding to each arm or RM $R_{k}$. We initialize the parameters for LinUCB by randomly initializing $b_{k}$ and setting parameter $A_{k}$ to the identity matrix. Based on the LinUCB algorithm, for each batch, the selected RM $R^{\star}_{t}$ is determined by $R^{\star}_{t}=R_{j}$ such that:
@@
where $\hat{\theta}_{k}=A_{k}^{-1}b_{k}$, and $\alpha$ is a parameter that controls the degree of exploration in LinUCB. A higher $\alpha$ encourages the algorithm to explore more aggressively by assigning higher uncertainty bonuses to actions with less information, while a lower $\alpha$ leads to more conservative behavior that prioritizes exploitation of known information. We provide an ablation study in Appendix˜D to examine the effect of $\alpha$ on the performance of LASeR. $A_{k}$ and $b_{k}$ are updated based on the MAB reward for each RM, which corresponds to the normalized negative train loss $\hat{\mathcal{L}}^{m}$ (described in detail in Sec.˜A.2):
@@
We emphasize that the effectiveness of our approach is not tied to particular bandit algorithm but rather from the exploration-exploitation trade-off of MAB framework. LASeR is agnostic to the bandit algorithms and can incorporate alternatives to LinUCB. Indeed, we provide a comparison of LinUCB and other bandit algorithms in Appendix˜C.
@@ 3.3 LLM and Bandit Training in LASeR
A key aspect of our approach is the generation of new preference training data in each iteration using the generations of the LLM itself and the RM selected by the MAB. Fig.˜1 presents our training procedure, broken down into three stages: (i) the MAB selects an RM $R^{\star}_{t}$ (see Sec.˜3.2; Fig.˜1 left), generating preference pairs by scoring the LLM’s outputs using the RM (Fig.˜1 (top-right)), and parameter updates to the LLM and MAB. In this way, the model continuously learns from its own outputs, guided by the selected reward model. After each LLM train step (i.e., one mini-batch), the MAB’s parameters are updated based on the observed MAB reward, i.e., how much the LLM’s loss decreased from using the selected RM. In the case of LinUCB, this involves updating the parameter estimates $b_{k},A_{k}$ (see Fig.˜1; bottom in green). This entire process – selection of reward models, generation of new supervision data, fine-tuning, and bandit updates – repeats for a total of $M$ iterations (summarized in Algorithm 1).
LASeR with Best-of-$n$ Sampling.  For settings where fine-tuning the LLM is not desirable or feasible, LASeR can also be applied to learn the MAB parameters without training the LLM. Rather than fine-tuning the model with preference data, we employ best-of-$n$ sampling (Lightman et al., 2023; Sun et al., 2024), where multiple responses are generated, and the best one is selected based on the RM. The bandit parameters are then updated using equation (2), with the MAB reward calculated as the negative normalized NLL loss on the train data. This updated bandit can subsequently be used for inference on the test set. This approach is particularly useful for long-context understanding tasks, where training would be too computationally intensive (see setting in Sec.˜4.2).
@@ 4 Experiments and Results
In this section, we evaluate LASeR across three domains: reasoning, instruction following, and long-context understanding. We compare our approach against baselines that utilize either a single RM or multiple RMs during LLM training. More detailed settings for both LASeR and the baselines are provided in Sec.˜A.1, and additional experimental results are presented in Appendix˜B.
@@ 4.1 Experimental Setup
Models.  We conduct our experiments on the Llama-3-8B base (AI@Meta, 2024) model. We present additional results with Mistral-7b-v3-Instruct (Jiang et al., 2023) and Qwen2.5-32B (Qwen Team, 2024a) in Appendix˜B. For training, all models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2021) for efficiency. For both training and inference, we do 0-shot prompting and sample n=30n\!=\!30 responses per prompt with temperature $0.8$ (see Sec.˜A.1 for more details).
Reward Models.  We select $K=4$ strong 7B RMs from RewardBench (Lambert et al., 2024), which include Zephyr-7B-Alpha, Qwen1.5-7B-Chat, Eurus-7B-KTO, and OLMo-7B-Instruct. Following the pipeline outlined in Lambert et al. (2024), for these models, we compute the reward for each response as the log-likelihood of the RM for that response (see Secs.˜A.1 and E for more details and discussion about the RM choices).
Datasets and Metrics.  Our experiments cover a range of tasks and datasets (see Sec.˜A.1):
@@
  • Long-Context Understanding: As fine-tuning LLMs on long-context inputs is computationally intensive, we demonstrate the effectiveness of LASeR using Best-of-$n$ sampling on LongBench (Bai et al., 2023), which consists of multiple tasks, such as single-document QA, multi-document QA, summarization, and few-shot learning. For summarization, we report Rouge-L (Lin, 2004), whereas for the rest, we report the F1 score.
@@
  • Best RM: From our collection of RMs, we pick the RM that corresponds to the best overall score on RewardBench (Lambert et al., 2024): Zephyr-7B-Alpha. We use this single RM during training (c.f. Sec.˜3.1). This baseline reflects the performance gain a user could expect when selecting the best RM from a leaderboard without knowing a priori how it generalizes to a particular task.
@@
  • Classifier Selection: To compare against a context-sensitive baseline that does not use an MAB, we train a $K$-way classifier to perform RM selection using data from RewardBench (see Sec.˜A.1). Specifically, for each query and RM, we compute the RM’s score of the annotated preferred and dispreferred response. The RM that assigns the correct preference ordering with the highest difference between the scores of the preferred and dispreferred responses is chosen as the RM label and used to train the classifier. While training the LLM, for each input $x_{i}\in\mathcal{D}$, we select the RM for building preference pairs based on this trained classifier.
@@
  • RM Agreement Ensemble: Because ensembling scores through averaging is sensitive to the absolute scores produced (which may differ between RMs), we follow Wang et al. (2024e) in ensembling through ranking and agreement. Specifically, we generate 32 responses for each query and sample 100 pairs for each. We score each pair with RMs, constructing a preference dataset by choosing the 10 pairs with the highest agreement of preference rankings across RMs.
@@
Conceptually, among the single RM selection methods, the best RM baseline serves as an “exploit-only” setting that exploits the best available RM based on aggregate RewardBench scores. On the other hand, the random and sequential selection baselines are “explore-only” in that they pick a new RM either randomly or via a predefined sequence, irrespective of the performance of each arm (RM). LASeR and the classifier approach represent two methods for selecting a single optimal RM for each instance; however, we note that the classifier approach depends on a fine-grained, in-distribution dataset of queries paired with their corresponding suitable RMs for annotating preferences, which is often absent in practice. In our experiments, we demonstrate the benefits of balancing the exploration-exploitation trade-off compared to "exploit-only" and "explore-only" approaches. Additionally, different RM ensemble baselines employ various methods for combining RMs while utilizing them simultaneously. Furthermore, we analyze potential conflicting signals among RMs to highlight the advantages of our framework over RM ensemble baselines (Sec.˜5).
@@ 4.2 Main Results
LASeR achieves the best average accuracy on reasoning tasks.  Table˜1 demonstrates that our method consistently outperforms the baselines across multiple reasoning benchmarks, particularly in the StrategyQA and GSM8K datasets. For example, LASeR improves by approximately $2\%$ absolute accuracy over the sequential baseline on both GSM8K and StrategyQA datasets.
@@
In cases where the best RM is not known beforehand, LASeR surpasses the performance of the average RM baseline by $2.7\%$ and the RM Score Ensemble baseline for each instance by $2.67\%$ (in accuracy averaged over the three datasets). Moreover, this lower performance by the RM Score baseline is not purely due to variance in the scores: LASeR also surpasses the RM Agreement Ensemble and RM Online Ensemble by $0.91\%$ and $2.27\%$, respectively. Overall, LASeR provides consistent results while the underlined second-place models show inconsistent performance across datasets. These results also emphasize the advantages of LASeR, as it eliminates the need to choose a different RM in advance or ensemble multiple RMs.
LASeR beats baselines at instruction-following.  Often, LLMs are used by large numbers of people with a diverse set of queries, goals, and intentions, and their preferences vary based on the underlying query. To demonstrate the effectiveness of LASeR in such settings, we compare the instruction-following performance in Fig.˜2, i.e., AlpacaEval win rates, of LLMs trained using LASeR with the baselines using WildChat. Specifically, LASeR achieves substantial win rates compared to single RM selection baselines such as random and sequential selection, with $78.33\%$, and $71.45\%$, respectively. We also outperform training with the single best RM (per RewardBench) by a $56.34\%$ win rate. Compared to RM ensemble baselines, LASeR achieves $72.69\%$ and $67.88\%$ win rates over RM Score Ensemble and RM Online Ensemble, respectively. We hypothesize the lower win rate of the baselines stems from the inability of these baselines to deal with conflicting signals from multiple RMs (see Fig.˜3 for further analysis). Overall, these results highlight that LASeR excels without gold labels and performs consistently well at following instructions across various user queries, showcasing its adaptability to diverse tasks.
LASeR’s adaptive RM selection helps long-context understanding.  Given the cost of training long-context systems, for LongBench (Bai et al., 2023), rather than fine-tuning a model using RMs, we employ the selected RM to rerank generation in Best-of-$n$ sampling (see Sec.˜3.3).
@@
In Table˜2, we observe that LASeR consistently outperforms the baselines across tasks. LASeR improves single-doc QA (SiQA) by $3.58$ F1 points over the base Llama-3-8B model and $2.64$ F1 points over random RM selection. On multi-doc QA (MuQA), our approach improves performance over the Llama-3-8B base model by $\approx\!4$ F1 points. Furthermore, on few-shot (FS) learning tasks, LASeR provides over $3$ points gain in F1 compared to the base model, surpassing the average RM performance by up to $2.4$ F1 points and demonstrating its effectiveness across tasks. Lastly, Table˜2 demonstrates that LASeR consistently outperforms the RM Score Ensemble baseline across different long-context tasks (except for the Summarization task where LASeR is comparable to Best RM), e.g., a $\approx\!3$ F1 point boost on single-doc QA and few-shot learning tasks.
@@ 5 Additional Analysis of LASeR
In this section, we present analyses on the presence of conflicts among RMs, which helps justify the benefits of LASeR, as well as its training efficiency compared to the baselines. We further provide additional analyses, including the generalization ability of LASeR in Appendix˜C, and ablation studies on LASeR’s design in Appendix˜D.
Presence of Conflicting Signals among RMs.  In Sec.˜4.2, we find that LASeR consistently outperforms other RM ensemble baselines across a wide variety of tasks. We attribute some of these performance gains to the inability of the multi-RM baseline to handle conflicting signals, resulting in subpar training data from multiple RMs. To study this, we sample pairs of outputs generated by Llama-3-8B on MMLU as well as WildChat and evaluate the consistency of response preferences measured by multiple RMs. Since pair-wise preferences are binary, we compute F1 to measure consistency with one RM’s preferences serving as the reference. Fig.˜3 (left on MMLU) reveals that Qwen and Zephyr have the highest agreement rate at $0.77$, while Qwen’s agreement with Eurus and Olmo is lower at $0.58$ and $0.43$, respectively. This is expected as Qwen and Zephyr are the top-performing models in reasoning according to RewardBench, while Olmo ranks the lowest in reasoning ability among the four models.
@@
We observe similar trends in agreement across RMs on WildChat (albeit with different agreement scores), which contains user queries asked LLMs in the wild; see Fig.˜3 (right). It appears that for more heterogeneous datasets with more categories, the level of disagreement among RMs (or conflict) increases. This also highlights LASeR’s advantages over multi-RM baselines that do not address conflicts in RMs and may explain why choosing one RM in LASeR and the best RM baseline outperforms multi-RM ensembles.
Training Efficiency of LASeR.
We show the accuracy-training time tradeoff in Fig.˜4 by comparing the GSM8K performance of training with LASeR and different baselines, along with the corresponding wall clock training time. Wall clock time is measured as the training time of a model (hours), keeping compute resource consistent. We find that sequentially optimizing over each RM performs the worst in terms of training time ($3\times$ of LASeR) while RM score ensemble has the worst accuracy (and takes $2\times$ the training time of LASeR). Moreover, LASeR outperforms all other baselines in terms of accuracy while maintaining the lowest training time,
@@
being more than twice as fast as the second-best baseline. LASeR’s efficiency comes from fast convergence compared to the sequential selection baseline and avoiding the overhead of loading and evaluating multiple RMs per step, as required in RM ensemble methods (Online Ensemble, Score Ensemble, and RM Agreement Ensemble). Instead, it selects and uses only one RM per mini-batch, saving GPU memory and compute. Ensemble baselines require scoring all candidate responses with all RMs, whereas LASeR scores responses with only the selected RM, reducing overhead. Lastly, we note that LASeR can be run in an offline data creation mode, where RM selection and scoring are done once to generate preference data, unlike online ensemble methods that require multiple RM passes per step.